At what temperature is the speed of sound in methane, CH4 equal to the speed of sound in air at room temperature?
1
Expert's answer
2020-10-01T10:09:48-0400

Expression for the speed of sound in methane:


"u=\sqrt{\frac{\gamma_1 RT_1}{\Mu_1}}"


The same speed of sound is in air (assuming that room temperature is 300 K):


"u=\sqrt{\frac{\gamma_2 RT_2}{\Mu_2}}"

Equate and solve for temperature of methane:


"\sqrt{\frac{\gamma_1 RT_1}{\Mu_1}}=\sqrt{\frac{\gamma_2 RT_2}{\Mu_2}},\\\space\\
T_1=T_2\frac{\Mu_1\gamma_2}{\Mu_2\gamma_1},\\\space\\
T_1=300\cdot\frac{16\cdot1.398}{29\cdot1.410}=183\text{ K},"

or -90.15°